OpenStudy (anonymous):

i will medal what is r twelfths = 1 and 9 tenths

OpenStudy (anonymous):

srry i put that on wrong

OpenStudy (texaschic101):

r/12 = 1 9/10 r/12 = 19/10 r = 19/10 * 12 r = 228/10 r = 22 4/5....that's 22 and 4/5

OpenStudy (anonymous):

tnx